OpenStudy (anonymous):

through: (-3,-3), parallel to y=7/3x+3

OpenStudy (jonnyvonny):

For, for something to be parallel, they have to have the same slope, or number, associated with the "x". So in this case, its "7/3". Now, we plug the slope and points into the equation "y=mx+b", to get the y-intercept: y=mx+b -3=(7/3)(-3)+b Wha'd you get? I'll post what I got